# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ VII Structural Signals of
The fact that there is a relation between reduced dynamics and the $n_{\bar{q}}=5.5$ data, inspired me to look for similar indicators of $\phi_{a}$ in the isoperimetric-quotient data for $R\lesssim 0.8$, where arrested dynamics appears to set in at nearly constant $\phi$. In this section, I therefore consider the properties of the $q$ distribution. I will also briefly comment on the inflection points to the dynamics that I introduced in Section V.
@@ VII.1 Large-Particle Pentagonal Neighborhoods
First, let us turn our attention to the large-particle-only distribution, i.e., the PDF that is obtained by only considering the $q$ values of the large-particle neighborhoods (the radical Voronoi diagram is not altered). This large-particle-only distribution is cut up into subdomains running from $q_{\mathrm{r}}(i-1)<q\leq q_{\mathrm{r}}(i)$ with $i=4$, $5$etc. On each of these subdomains, I establish the (sub)mode of the distribution — the value of $\phi$ for which the highest (local), large-particle-only PDF is found. Here, I specifically examined pentagonal and hexagonal neighborhoods, as these are most prevalent in the bidisperse disk system at high values of $\phi$, also see Appendix B. The value of the PDF associated with this submode, denoted $P^{\ast}_{i}$, is shown in Fig. 5a-c as a function of $\phi$ for the two domains $i=5$ (blue; square to pentagon) and $i=6$ (red; pentagon to hexagon), respectively.
@@
Figure 5: Identifying the arrested dynamics via the isoperimetric-quotient distribution. (a-c) The (natural) logarithm of the isoperimetric-quotient PDF for large particles only, evaluated at its domain-restricted mode value $P^{\ast}_{i}$, as a function of the area fraction $\phi$. The main text details the procedure by which this data was obtained. Blue points show the value for the submode restricted to the domain $q_{\mathrm{r}}(4)<q\leq q_{\mathrm{r}}(5)$ ($P^{\ast}_{5}$) and red points are obtained from data restricted to $q_{\mathrm{r}}(5)<q\leq q_{\mathrm{r}}(6)$ ($P^{\ast}_{6}$), respectively. The vertical dashed purple line indicates the local minimum, in this representation this is the location of the cusp in $P^{\ast}_{5}$, while the green one indicates the $\phi$, for which the red and blue curves cross each other for the first time. In all cases, the value of the size ratio is labelled in the bottom-left corner of the panel. The error bars indicate the standard error of the mean, though in panel (a) there is a clear set of outliers. These were ignored in my analysis. (d) Reproduction of part of the $R$-$\phi$ state diagram of Fig. 4. The orange and magenta data shows $\phi$ associated with the minimum in the large-particle $P^{\ast}_{5}$ and the cyan data indicates the inflection point in the diffusion data, see Section V. The black point labelled “a” corresponds to the glass transition found in Ref. [1]. The black point labelled “b” represents the “onset glass transition” as reported in Ref. [2]. All thick, light-colored curves serve as guides to the eye.
@@
Note that $P^{\ast}_{5}$ has a clear minimum, see Fig. 5a-c. Taking the natural logarithm of $P^{\ast}_{5}$ gave this minimum the appearance of a cusp. Both sides of this cusp were subsequently fitted using polynomials and the intersection point computed  [^6] Note that $P^{\ast}_{5}$ is a positive number always, hence the cusp is not a sign of a number going through zero, as is often found in logarithmic representations where the absolute of a value is plotted.. The associated value of $\phi$, say $\phi^{\ast}$, is shown in Fig. 5d, where the error bar gives an indication of the accuracy of the fit. The value of $\phi^{\ast}$ for $R>0.83$ is indicated in magenta and for $R<0.83$ in orange, respectively. I will return to the need for this distinction shortly. I also examined the distribution $P_{\phi}(q)$, which derives from all particle neighborhoods, and the PDF that results by considering only those neighborhoods that belong to small particles. In both distributions, the effect for the square-to-pentagon subdomain was clearly present for values of $R\uparrow 1$, but for $R\lesssim 0.83$ there were only small hints of a feature in the distribution, which proved difficult to fit, see the ESI [59]. This further justifies my choice for examining only the large particles.
Large particles, especially for $R\ll 1$, should be the least mobile and, thus, arrested dynamics should originate with these particles. This could explain why I see the strongest signal originating from this particle fraction. I find that the $q$ associated with $P^{\ast}_{5}$ shifts ever closer to $q_{\mathrm{r}}(5)$ as $\phi$ is increased, with the peak-value first increasing, then decreasing rapidly, see the ESI for additional details [59]. In all cases, the decrease sets in as the large-particle neighborhoods tend toward hexagonal, i.e., beyond the large-particle crossover $P^{\ast}_{5}=P^{\ast}_{6}$. This is when the large particles start to form hexagonal-like clusters to allow the system to further densify.
@@
Note that whenever the system has a sufficiently large size asymmetry, the presence of the small particles interferes with the clustering and $P^{\ast}_{6}$ reduces as $\phi$ is increased further. Following the vertical purple dashed lines, there is a strong correlation between this downward trend in $P^{\ast}_{6}$ and the presence of the minimum in $P^{\ast}_{5}$ (for sufficiently small $R$). However, the maximum in $P^{\ast}_{6}$ proved more difficult to fit and I therefore did not do this. The intuition is that if the dynamics of the system drastically changes, one would expect there to be frustration. This will lead to local neighborhoods distorting, e.g., see Ref. [54], rather than the system favoring neighborhoods that are close to $q_{\mathrm{r}}(n)$. This can also be seen in Fig. 2a-c.
Examining the orange data in Fig. 5d, I note that the extracted value of $\phi^{\ast}$ lies somewhere between the curves $\max_{\phi}\!\left(n_{\bar{q}}\right)$ and $\phi_{a}$. Thus, within the limitations of my numerical analysis, I do not find a direct structural indicator of ‘arrested dynamics’, but one that is very close to it. I provide an additional analysis in Appendix E that is moderatly succesful in establishing the $\phi_{a}$ line. This suggests that slightly more extensive local measures of the structure, say also accounting for the next-nearest neighbors, would likely be able to locate $\phi_{a}$ more accurately.
Intriguingly, the minima in $P^{\ast}_{5}$ are also present for $R\gtrsim 0.83$, as indicated by the magenta data in Fig. 5d. It appears that the mean value of $\phi^{\ast}$ on this domain is slightly lower than that for smaller $R$ (orange data). That is, $\phi^{\ast}$ lies closer to the green dashed line. This shows, to the best of my understanding for the first time, a structural indicator of the system changing within the crystalline phase. Though it should be noted that what I am picking up on is a very small effect. The same effect is present when only focusing on $P_{\phi}(q_{\mathrm{r}}(5))$. Examination of system snapshots for $\phi\approx 0.777$ and $R\gtrsim 0.83$ does not reveal anything profound in the structure of the crystalline phase. That is, the particle arrangements appear to follow a standard hexagonal pattern, without obvious signs of defects or dislocations. This is, however, expected as the presence of pentagonal neighborhoods has dropped off to nearly zero, see Appendix B. The signal that $P^{\ast}_{5}$ picks up is thus a signpost, the origin of which requires interpretation, which I will provide in Section VIII.
@@ VII.2 Inflection Point for the Diffusivity
Figure 5d also shows the inflection point data (cyan) that was obtained from the measurement of the diffusion coefficient, see Section V. This could only be accurately established for $R\lesssim 0.83$, hence it is not shown over the entire range. There is a substantial error in the data, because it is a difficult point to extract. Note, however, that within this error, the associated value of $\phi$ is effectively constant. This is similar to the way $\phi_{a}$ behaves on this domain. Unlike for $\phi_{a}$, I was unable to determine a structural feature that I could convincingly correlate to the $\phi$ of the inflection point.
I suspect that this inflection point is set by the dynamics of the system changing from freely diffusing in a medium-dense environment to cage-like. The caging transition could not be readily obtained from the data that was available in Lozano et al. [1]. However, I note that Li et al. [2] studied a 55:45 (close to the 50:50 system) of small and large spheres (2.08 $\mu$m and 2.91 $\mu$m, respectively, resulting in $R\approx 0.72$) in quasi-2D confinement. I will thus rely on this data to make a comparison, see the black point labeled “b” in Fig. 5d, under the assumption that the small change in stoichiometry does not affect the system significantly.
The authors of Ref. [2] perturb their sample using a focused laser and track the particle displacement that this causes. They report a peak in the associated displacement values around $\phi\approx 0.6$, which I have used in my comparison. This peak is collocated with the onset of glassy dynamics according to the authors. However, despite the apparent similarities to the peak in Ref. [1] the origin seems different. That is, the MSD data in Ref. [2] does not reveal a clear glass transition; if such a transition is present, it may be that it occurs above $\phi>0.73$. Thus, Li et al. interpret their point as being related to the formation of cages and therefore dub it an “onset glass transition.”
Despite the differences between Refs. [1] and [2], the indication seems that the signals picked up on by the respective approaches have to do with the dynamics of the local neighborhoods. The presence of a structural indicator appears more fortuitous than profound, but I will show that this is not the case next.
@@ VIII Geometric Understanding
The above evidence points to the significance of the area fraction $\phi\approx 0.777$. In line with previous studies into this topic, the structural signifiers of (apparent) arrested dynamics are relatively weak [54, 53]. My shell-percolation approach in Appendix D tracks $\phi_{a}$ very accurately. However, it relies on a large amount of data to obtain a result of this quality, as well as on comparison between closely spaced values of $\phi$ to arrive at an appropriate match to $\phi_{a}$. Similarly, the identification which is based on the cusp in $\max_{\phi}\!\left(n_{\bar{q}}\right)$, see Section VI, and the cusp in the large-particle-only abundance of pentagonal neighborhoods, see Section VII.1, require large data sets to identify a small, but real effect. I also considered a form of Voronoi entropy [92] based on the isoperimetric quotient, which showed a weak signal near $\phi_{a}$; it was therefore not reported on here.
Nonetheless, a structural effect is present and I will thus explore this aspect further. The ‘ideal’ or ‘pure’ structural signal, I must conclude, is somehow obscured. To uncover it, I turn to a granocentric-inspired [93, 94, 95] view of the particles. This choice is also informed by the fact that I make a tentative connection to random loose packing in Appendix A. That is, the granocentric approach has been applied to study random packings. The results in Section VII.1 show that it is sensible to focus on $R=1$, as there are features of this system that are shared by those with $R\lesssim 0.8$ that do not crystallize. I will now argue how using a granocentric approach leads to a structure-based transition at $\phi_{g}=\sqrt{3}\pi/7\approx 0.777343$.
VIII.1 Introducing the Floret Pentagonal Tiling
@@
It is straightforward to show that a small particle can be stabilized by being surrounded by 5 larger particles, leading to a pentagonal neighborhood. At contact, this occurs for $R\approx 0.701302$; the ESI [59] provides the analytic expression. Large particles can have a perfectly heptagonal neighborhood, whenever they are surrounded by 7 small particles. At contact, this occurs for $R\approx 0.766422$, also see the ESI [59]. Neither of these values corresponds particularly well with the transitions observed in the system. It should be clear that the presence of such neighborhoods is a rarity. Thus, I would not expect the appearance of arrest to be captured by such a simple picture. Interference with crystallization should ideally be shared equally among all particles.
@@
Figure 6a shows a single, pentagonal neighborhood of a central disk surrounded by five neighbors, which meets this requirement. It is easy to demonstrate, see the ESI [59], that the volume fraction associated with this central particle is $\phi=\sqrt{3}\pi/7$. These neighborhoods can be assembled into a regular structure that is effectively a close-packed hexagonal crystal with one particle out of every 7 missing, see Fig. 6b. For each collection of 6 particles surrounding a void, henceforth referred to as ‘wheel’, every particle is in contact with 5 neighbors. Thus, every associated Voronoi cell has exactly 5 vertices. This also means that the particle is completely immobilized, independent of whether it is frictional or non-frictional.
@@
Figure 6: Geometric ground state that can be related to the glass transition. (a) The Voronoi cage (purple) around a single particle (red) surrounded by 5 neighboring disks (blue). These particles are in contact (green points). The black dots indicate the centers of the respective disks. (b) Organizing the particles into ‘wheel’-like neighborhoods, wherein all the Voronoi cells have five vertices. The central wheel (blue) shows the 6 pentagonal Voronoi cells (purple). The six surrounding wheels show that these can be stacked into a crystalline structure. The use of alternating red and green coloring is meant to help distinguish the wheels, but otherwise carries no physical significance.

### v2

@@ VII Structural Signals of
The fact that there is a relation between reduced dynamics and the $n_{\bar{q}}=5.5$ data for $R\gtrsim 0.83$, inspired me to look for similar indicators of $\phi_{a}$ in the isoperimetric-quotient data for $R\lesssim 0.8$, where arrested dynamics appears to set in at nearly constant $\phi$. Here, I therefore consider the properties of the $q$ distribution and comment on the inflection points obtained in Section V.
@@ VII.1 Large-Particle Pentagonal Neighborhoods
First, consider the large-particle-only distribution, i.e., the PDF that is obtained by only considering the $q$ values of the large-particle neighborhoods (the radical Voronoi diagram is not altered); this will be justified shortly. This large-particle-only distribution is cut up into subdomains running from $q_{\mathrm{r}}(i-1)<q\leq q_{\mathrm{r}}(i)$ with $i=4$, 55etc. On each of these subdomains, I establish the (sub)mode of the distribution — the value of $\phi$ for which the highest (local), large-particle-only PDF is found. Examining the peak values amplifies subtle changes in the underlying distributions, though one should be careful not to overinterpret the outcome. The goal of this section is solely to show that extrapolating $\phi\approx 0.777$ to $R=1$ is justifiable, even if the effects that are picked up on are small. Here, I specifically examined pentagonal and hexagonal neighborhoods, as these are most prevalent in the bidisperse disk system at high values of $\phi$, also see Appendix B. The value of the PDF associated with this submode, denoted $P^{\ast}_{i}$, is shown in Fig. 7a-c as a function of $\phi$ for the two domains $i=5$ (blue; square to pentagon) and $i=6$ (red; pentagon to hexagon), respectively.
@@
Figure 7: Identifying the arrested dynamics via the isoperimetric-quotient distribution. (a-c) The (natural) logarithm of the isoperimetric-quotient PDF for large particles only, evaluated at its domain-restricted mode value $P^{\ast}_{i}$, as a function of the area fraction $\phi$. The main text details the procedure by which this data was obtained. Blue points show the value for the submode restricted to the domain $q_{\mathrm{r}}(4)<q\leq q_{\mathrm{r}}(5)$ ($P^{\ast}_{5}$) and red points are obtained from data restricted to $q_{\mathrm{r}}(5)<q\leq q_{\mathrm{r}}(6)$ ($P^{\ast}_{6}$), respectively. The vertical dashed purple line indicates the local minimum; in this representation this is the location of the cusp in $P^{\ast}_{5}$. The vertical dashed green line indicates the $\phi$ for which the red and blue curves cross each other for the first time. In all cases, the value of the size ratio is labelled in the bottom-left corner of the panel. The error bars indicate the standard error of the mean, though in panel (a) there is a clear set of outliers, caused by limited equilibriation and subsequent issues in the fitting. These were ignored in my analysis. (d) Reproduction of part of the $R$-$\phi$ state diagram of Fig. 6. The orange and magenta data shows $\phi$ associated with the minimum in the large-particle $P^{\ast}_{5}$ and the cyan data indicates the inflection point in the diffusion data, see Section V. The black point labelled “a” corresponds to the glass transition found in Ref. [1] and the black bar labelled “b” indicates the range for the jamming transition from Refs. [2, 3]. The black point labelled “c” represents the “onset glass transition” as reported in Ref. [4]. All thick, light-colored curves serve as guides to the eye.
@@
Note that $P^{\ast}_{5}$ has a clear minimum, see Fig. 7a-c. Taking the natural logarithm of $P^{\ast}_{5}$ gives this minimum the appearance of a cusp. Both sides of this cusp were subsequently fitted using polynomials and the intersection point computed  [^11] Note that $P^{\ast}_{5}$ is a positive number always, hence the cusp is not a sign of a number going through zero, as is often found in logarithmic representations where the absolute of a value is plotted.. The associated value of $\phi$, say $\phi^{\ast}$, is shown in Fig. 7d, where the error bar gives an indication of the accuracy of the fit. The value of $\phi^{\ast}$ for $R>0.83$ is indicated in magenta and for $R<0.83$ in orange, respectively. The orange data appears to follow the $\max_{\phi}\!\left(n_{\bar{q}}\right)$ curve, albeit at a slightly lower value of $\phi$, which hints at a correlation betweent the two measurements. For $R>0.83$ the data appears nearly constant and closer to the $\phi=0.777$ line. I used this indicator to justify extending the $\phi=0.777$ line to $R=1$.
I also examined the distribution $P_{\phi}(q)$, which derives from all particle neighborhoods, and the PDF that results by considering only those neighborhoods that belong to small particles. In both distributions, the effect for the square-to-pentagon subdomain was clearly present for values of $R\uparrow 1$, but for $R\lesssim 0.83$ there were only small hints of a feature in the distribution, which proved difficult to fit, see the ESI [64]. This further justifies my choice for examining only the large particles. Moreover, large particles, especially for $R\ll 1$, should be the least mobile and, thus, arrested dynamics should originate with these particles. This could explain why I see the strongest signal originating from this fraction.
@@
The data shows that the $q$ associated with $P^{\ast}_{5}$ shifts ever closer to $q_{\mathrm{r}}(5)$ as $\phi$ is increased, i.e., the neighborhoods become increasingly pentagonal. The peak-value first increases, then decreases rapidly, see the ESI for additional details [64]. In all cases, the decrease sets in as the large-particle neighborhoods tend toward hexagonal, i.e., beyond the large-particle crossover $P^{\ast}_{5}=P^{\ast}_{6}$. This is when the large particles start to form hexagonal-like clusters to allow the system to further densify. Note that whenever the system has a sufficiently large size asymmetry, the presence of the small particles interferes with the clustering and $P^{\ast}_{6}$ reduces as $\phi$ is increased further, as local neighborhoods distort [58]. Following the vertical purple dashed lines in Fig. 7a-c, there is a strong correlation between this downward trend in $P^{\ast}_{6}$ and the presence of the minimum in $P^{\ast}_{5}$ (for sufficiently small $R$). However, the maximum in $P^{\ast}_{6}$ proved more difficult to fit, so that I could not ascertain the degree of correlation.
Intriguingly, the minima in $P^{\ast}_{5}$ are also present for $R\gtrsim 0.83$, as indicated by the magenta data in Fig. 7d. It appears that the mean value of $\phi^{\ast}$ on this domain is slightly lower than that for smaller $R$ (orange data). That is, $\phi^{\ast}$ lies closer to the orange dashed line. This shows, to the best of my understanding, a structural indicator of the system changing within the crystalline phase. However, it should be noted that the effect is difficult to obtain and quite subtle.
@@ VII.2 Inflection Point for the Diffusivity
Figure 7d also shows the inflection point data (cyan) that was obtained from the measurement of the diffusion coefficient, see Section V. This could only be accurately established for $R\lesssim 0.83$, hence it is not shown over the entire range. There is a substantial error in the data, because it is a difficult point to extract. Note, however, that within this error, the associated value of $\phi$ is effectively constant. This is similar to the way $\phi_{a}$ behaves on this part of the $R$ domain. The guide to the eye in Fig. 7d is a fitted constant with value $\phi\approx 0.607$ to all the data points. This value lies close to where Li et al. [4] locate their onset glass transition or equivalently the appearance of caging in the fluid, see the black point labeled “c” in Fig. 7d. This is added here under the assumption that the small change in stoichiometry in Ref. [4] (0.55:0.45 rather than 1:1) does not affect the system significantly. Thus, the appearance of the inflection point can be given a caging interpretation, meaning that a new pathway for $D$ reduction becomes available. Note that the caging transition could not be readily obtained from the data that was available in Lozano et al. [1], as was discussed earlier in the introduction.
@@ VIII Geometric Understanding
Turning to the literature on foams and tissues [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95], I consider it possible (and plausible) that a geometric feature underlies the observations of dynamic arrest in the disk systems. This would be analogous to the way the pentagonal shape controls the onset of jamming in model tissues. To identify this feature in the disk systems, I adopt a granocentric view [101, 102, 103]. This choice is also informed by the apparent connection to frictional jamming and RLP, for which the method was originally developed. Following this approach, I studied a single disk surrounded by a single neighbor shell of disks.
@@
It proved difficult to explain features of the state diagram as a function of $R$ using this approach on a bidisperse system. That is, it is straightforward to show that a small particle can be stabilized by being surrounded by 5 larger particles (all in contact), leading to a pentagonal neighborhood. At contact, this occurs for $R\approx 0.701302$; the ESI [64] provides the analytic expression. Large particles can have a perfectly heptagonal neighborhood, whenever they are surrounded by 7 small particles (all in contact). At contact, this occurs for $R\approx 0.766422$, also see the ESI [64]. Neither of these values corresponds particularly well with the ‘critical point’ to the crystalline region (as a function of $R$), which occurs for $R\approx 0.83$. It should be clear that the presence of such neighborhoods is a rarity. Thus, I would not expect the appearance of arrest to be captured well by such an analysis. However, the granocentric approach did prove beneficial in making the connection to Voronoi-based foam and tissue models, when applied to a monodisperse system.
The above evidence points to the significance of the area fraction $\phi\approx 0.777$ (and potentially $\approx 0.607$). It is reasonable to assume that the aspect ratio $R$ is not significant in setting the transition criterion, given that $\phi_{a}$ is nearly constant over a wide range of $R$. Note that here it is important that this holds for amorphous packings, and is thus different from the appearance of constant $\phi$ lines in (quasi-)crystalline packings [125]. Using the extrapolation to $R\uparrow 1$ as I argued for above, it becomes possible to examine the properties of a monodisperse system only for clues to the origin of this value of $\phi$. Additionally, the experimental results for a wide range of length scales [82, 1, 3], appear to indicate that the particle size $\sigma$ and temperature $T$ do not control the transition. To help support the former insight, I make the connection to the literature on experiments and simulation of frictional systems in Appendix A.
VIII.1 The Floret Pentagonal Tiling
Considering instead hexagonal neighborhoods of equal-sized disks, reveals the following. Figure 8a shows a single, pentagonal neighborhood of a central disk surrounded by five neighbors. This configuration can be seen as a limiting state of moving a sixth neigboring disk upward until it is no longer a Voronoi neighbor with the central particle. It is easy to demonstrate, see the ESI [64], that the area fraction associated with this central particle is $\phi=\sqrt{3}\pi/7\approx 0.777343$, which is remarkably close to the value of $\phi$ for which dynamical arrest was observed in the event-driven simulations.
@@ VIII.1 The Floret Pentagonal Tiling
Interference with hexagonal crystallization — leading to amorphous states with dynamic arrest — should ideally be shared equally among all particles. The pentagonal neighborhood identified in Fig. 8a can be assembled into a regular structure that is effectively a close-packed hexagonal crystal with one particle out of every 7 missing, see Fig. 8b. For each collection of 6 particles surrounding a void, henceforth referred to as “wheel”, every particle is in contact with 5 neighbors. Thus, every associated Voronoi cell has exactly 5 vertices. This also means that each particle is completely immobilized, independent of whether it is frictional or non-frictional. The area fraction of this configuration is $\phi=\sqrt{3}\pi$ and it is identical to the isoperimetric ratio $q$.
@@
Figure 8: Geometric ground state that can be related to the glass transition. (a) The Voronoi cage (purple) around a single particle (red) surrounded by 5 neighboring disks (blue). These particles are in contact (green points). The black dots indicate the centers of the respective disks. (b) Organizing the particles into wheel-like neighborhoods, wherein all the Voronoi cells have five vertices. The central wheel (blue) shows the 6 pentagonal Voronoi cells (purple). The six surrounding wheels show that these can be stacked into a crystalline structure. The use of alternating red and green coloring is meant to help distinguish the wheels, but otherwise carries no physical significance.